### v1

@@ 4.1 Planets around Sun-like Stars
Focusing initially on planets orbiting Sun-like stars, which for our purposes are stars within ±5% of the Sun’s mass, we observe distinctive patterns in occurrence rates across different planet categories. The occurrence rate for a given planet type is defined as
For the run, the occurrence rates in the different components are shown in Figure 6. The planet populations in the bulge and thin disc are remarkably similar, the same is true for the thick disc and halo. This similarity is primarily attributed to these components covering similar metallicity regimes, which significantly influences planet occurrence rates in our model. A list of the the occurrence rates can be found in Table 7.
@@
The thick disc and halo bear a distinct planet population compared to the more metal-rich regions. The lower metallicity of the stars favor low-mass planet formation over more massive ones. The occurrence rate of Earth-like planets is a 1.5 times larger than for the bulge and thin disc component, at 5.5 planets per star, with a comparable drop in the occurrence rate of super-Earths to 2.5 planets per star. More striking is the difference in occurrence rates for Neptunian and giant planets. Neptunian planets are 5 times more common in the thin disc and bulge compared to the thick disc and halo, with frequencies around 0.6. This ratio increases with increasing planet mass, the occurrence rates of giant planets with masses are 20 times higher in the metal-rich components compared to the thick disc and halo. The planets in these regions are also older on average, at 7.7 Gyr for the thick disc and 8.2 Gyr for the halo (see also Table 5).
@@
Taking a higher level view, we analysed the distribution of planet types across the whole galaxy. We show radial profiles in Figure 7, while 2D distribution maps are shown in Figure 8. Giant planets have a higher concentration in the center of the galaxy, correlating with the region’s higher metallicity. Figure 7 shows the normalised planet density as a function of distance from the galactic center. The spatial density of planets depends on the stellar density and planet occurrence rates. Thus, unsurprisingly, the highest density of planets is found towards the galactic center due to the large stellar density in this region. Moving away from the galactic center, the number of giant planets drops faster than for lower mass planets, due to the higher metallicity-dependence of the giant planet occurrence rates. In total, 90% of giant planets and 84% of Earth-like planets can be found within 8 kpc from the galactic center, roughly the distance of the Sun from the Milky Way center.
Bulge |
Thin Disc |
Thick Disc |
Halo |
Earth-Like |
8.0 |
6.1 |
7.8 |
8.3 |
Super-Earth |
7.5 |
5.7 |
7.5 |
8.0 |
Neptunian |
6.8 |
5.1 |
7.2 |
7.3 |
Sub-Giant |
6.9 |
5.1 |
6.9 |
6.6 |
Giant |
6.5 |
4.8 |
6.8 |
6.1 |
@@
Further, we find that the total number of planets is dependent on the assumed number of initial planet embryos. While the dependence is relatively weak for massive planets - echoing the conclusions of Emsenhuber et al. (2021a) - a more pronounced dependence is observed for lower mass planets. We calculate the total number of planets within kpc of the HESTIA galaxy for = 10, 20, 50 an 100 and tabulate them in Table 6 . Depending on the assumed number of embryos, this yields between and planets around Sun-like stars in the HESTIA Analogue. It is probable that the number of embryos is not universal for protoplanetary systems and depends, amongst other things on the metallicity (see Section 5.3.2). The = 10 and = 100 cases can however be treated as a lower and upper estimate for the total number of planets.
10 |
20 |
50 |
100 |
Earth |
3.05 |
5.18 |
9.36 |
11.00 |
Super-Earth |
3.73 |
5.14 |
9.09 |
10.97 |
Neptunian |
1.25 |
1.37 |
1.31 |
1.42 |
Sub-Giant |
0.35 |
0.38 |
0.29 |
0.32 |
Giant |
0.82 |
0.82 |
0.71 |
0.74 |
Total |
9.21 |
12.89 |
20.77 |
24.45 |
@@ 4.2 Low-Metallicity Environments
To quantify this effect, we consider an alternative metallicity model, where we assign zero planets to stars below the metallicity threshold [Fe/H] = -0.6, and compare occurrence rates to the previous prescription. Figure 9 shows the ratio between the occurrence rate for the fiducial and this truncated model for the different planet types and galaxy components. The variations are minimal for the bulge and thin disc, since most stars in these regions are above the metallicity threshold. For the low-metallicity environments, the occurrence rates of low-mass planets in the halo is up to 50% lower when assuming this more restrictive metallicity assumption.
Interestingly, the occurrence rate for giant planets does not significantly differ between the two models, not even in the low-metallicity galaxy components like the halo, even though the occurrence rates of these planets have the highest metallicity dependence. This stems from the fact that, within our model, giant planets already show near-zero occurrence rates at a metallicity of [Fe/H] = -0.6. Consequently, the variation in the metallicity models does not significantly change their estimated frequency. In contrast, the occurrence rates for Earth-like planets show severe reduction. Given that the occurrence of lower mass planets is only weakly dependent on metallicity, and considering that approximately half of the star particles in the halo have metallicities below our cutoff, the occurrence rate for Earth-like planets is effectively halved in this scenario.
It is important to note that the actual Milky Way halo is more metal-poor than that of the HESTIA analogue. The effect demonstrated here is therefore likely to be even more pronounced in reality.
@@ 4.3 Planets around Intermediate- and Low-Mass Dwarfs
For lower mass stars, the correlation between planet occurrence rates with metallicity becomes weaker, although the average metallicity to form a given planet type is larger (Burn et al., 2021). Consequently, the planet populations in the different regions of the galaxy are more homogeneous compared to those around Sun-Like stars. The occurrence rates are shown in Figure 10 and Table 7.
Around intermediate mass dwarfs (), giant planets are extremely rare in the thick disc and halo with occurrence rates of 0.001 – 0.002. These values increase by a factor of 10 for the metal-rich regions. The occurrence rate of Earth-like planets for these stars is rather insensitive to the metallicity and and stays between 4 and 5 planets per star. Super-Earths on the other hand are more metallicity-dependent with occurrence rates around 1 planet per star for the metal poor region, three times lower than for the metal rich regions.
@@ 5.1 Planet Populations in the Galactic Context
Our results show that galactic chemical evolution and variations in stellar populations profoundly affect planet demographics throughout the galaxy. For Sun-like stars, we observe that metal-rich regions, specifically the bulge and thin disc of the Milky Way, harbor the largest number of planets with masses , in both absolute numbers and relative frequencies. In contrast, the metal-poor regions, i.e. the thick disc and halo, show a significant suppression in giant planet formation, which is attributed to the lack of solid material building blocks, which are crucial for forming the massive planet cores before the dispersal of the protoplanetary gas discs. Specifically, the occurrence rate of giant planets around Sun-like stars is an order of magnitude higher in the metal-rich regions. Earth-like planets and Super-Earths are more insensitive to the metallicity variations, but still affected. Earth-like planets around Sun-like stars are actually most frequent in the low metallicity environments of the thick disc and halo, being around twice as common as in the thin disc. Super-Earths show the opposite trend and are about half as frequent in the low metallicity regions.
For lower-mass stars, with and , we observe that planet populations are less dependent on metallicity, leading to a more uniform distribution across the galaxy. The lower amounts of solid material available in the protoplanetary discs around these stars generally results in host stars having a higher average metallicity compared to the Sun-like case for the same planet type. Thus, giant and sub-giant planets remain significantly rarer in low-metallicity regions. No giant planets with masses greater are found around any stars of masses , regardless of the galactic environment.
Occasionally, the concept of a galactic habitable zone is discussed in the literature (Gonzalez et al., 2001; Lineweaver, 2001; Prantzos, 2008; Gowanlock et al., 2011), which suggests that life-supporting conditions are more likely in certain galactic regions, including some of our previous work involving hydrodynamical galaxy simulations (Forgan et al., 2017). The key idea is that the galactic center, with its extreme environmental conditions, is less conducive to life due to higher radiation levels and more frequent stellar encounters. Conversely, the outer regions are thought to be too metal-depleted for habitable planet formation. While we have not explored the radiation-related arguments in the inner region, we find no evidence of an outer limit. Low-mass, rocky planets remain abundant throughout the galaxy far into the stellar halo. Moreover, some of these models suggest that giant planets might be detrimental to habitability or the formation of Earth-like planets (Lineweaver, 2001; Forgan et al., 2017). Under this assumption, stars in the far-out metal-poor regions could be considered favourable, since giant planet formation is suppressed in these regions.
@@ 5.2 Empirical Occurrence Rates
The robustness of our results applied to the real Milky Way depends on the accuracy of the NGPPS occurrence rates compared to true occurrence rates. Matuszewski et al. (2023) conducted a planet yield analysis for the PLATO mission (Rauer et al., 2014), and compared the occurrence rate estimates around FGK stars from various sources. This comparison includes the NGPPS run, and two empirical models based on Kepler data - one by Hsu et al. (2019) and the other by Kunimoto & Matthews (2020). Unlike the mass-based categorization in NGPPS, these models focus on planet radius, aligning with the observational constraints of transit missions.
@@ 5.3.1 Disc Lifetimes
The lifetime of protoplanetary discs, primarily governed by external photoevaporation, plays a crucial role in planet formation. We adopt the range of photoevaporation rates from the Emsenhuber et al. (2021b) NGPPS sample, which are chosen to reproduce typical disc lifetimes in the solar neighbourhood. Efficient massive planet formation () in the Bern model require disc lifetimes around 3 – 6 Myr, meaning early disc dispersal is not a dominant mechanism to halt giant formation (Schlecker et al., 2021), which is reflected in the weak correlation coefficient between photoevaporation rate and planet type (Figure 2).
However, it is important to note that most studies of disc lifetimes, including those by Emsenhuber et al. (2021a) and Ansdell et al. (2017), have focused on regions with relatively low stellar densities. In contrast, the galactic center presents a drastically different environment, where higher stellar densities, lead to more intense photoevaporation and consequently, shorter disc lifetimes (Winter & Haworth, 2022). The Central Molecular Zone (CMZ) in particular, exhibits gas surface density of (Henshaw et al., 2016), exceeding that of the solar neighbourhood by a factor of 100 (Kruijssen, 2012), which is connected to larger and denser stellar clusters (Farias et al., 2023). This heightened density can significantly amplify photoevaporation, and thus reduce disc lifetime. Winter et al. (2020) constructed an analytical model for the disc lifetime as a function of galactic environment, and find that virtually all discs in the CMZ have lifetimes ¡3 Myr, severely hindering giant planet formation.
While the CMZ is an exceptional region within the Milky Way, with a diameter of approximately 0.6 kpc, it is not sufficiently resolved in our simulations. Nonetheless, it is plausible that the shortened disc lifetimes in high-density regions towards the galactic center could lower the occurrence rates of giant planets.
@@ 5.3.2 Initial Embryo Distribution
The formation of planet embryos is not explicitly modeled in the Bern model. Instead, a fixed number of embryos are distributed logarithmically at the beginning of each run. However, an analysis of the connection between initial parameter and resulting planet population for the NGPPS population performed by Schlecker et al. (2021) find the location of a planet embryo is the single strongest predictor for the final planetary type, and this distribution might vary significantly across different galactic environments.
@@
In terms of the number of planet embryos, they find that the total number of embryos increases with increasing planetesimal disc mass and density slope. For low solid disc masses, specifically between 0.5 and 5 AU, around 20–40 embryos are formed, depending on the assumed density profile. At a higher planetesimal disc mass of , around 60 embryos are formed, irrespective of the assumed profile. Translated into the context of our planet population model, which assumes universal distribution of gas disc masses the galaxy, these findings imply that stars in metal-rich regions –boasting higher solid disc masses– would, on average, form a greater number of planet embryos. Since the occurrence rates of low-mass planets in the model increases with the initial number of embryos (Table 6), this would indicate that Earth-like planets and Super-Earths are more common in metal-rich regions. The frequencies of giant planets on the other hand are insensitive to and are not strongly affected.
5.4 Relation to Previous Work
@@
Bashi & Zucker (2022) analysed a sample of 506 planet candidates orbiting around 369 FGK host stars, categorising them into thin and thick disc components. They found that the occurrence rate of close-in super-Earths in the thin disc is approximately 1.3-3.6 larger than in the thick disc, albeit with low number statistics as only 8 of their planets are associated with thick disc stars. This is consistent with our findings, where the lower average [Fe/H] in the thick disc leads to an increased occurrence rate of Earth-like planets, with a corresponding drop in the frequency of super-Earths by a factor of 1.5 . Bashi & Zucker (2022), however, attribute this variation primarily to the dynamical history of the stars, including potential disruptive encounters, and gravitational perturbations among planets in typically older systems, rather than differences in metallicity.

### v2

@@ 4.1 Planets around Sun-like stars
Focussing initially on planets orbiting Sun-like stars, which (for our purposes) are stars within ±5% of the Sun’s mass, we observe distinctive patterns in occurrence rates across different planet categories. For the run, the occurrence rates in the different components are shown in Figure 6. The planet populations in the bulge and thin disc are remarkably similar, the same is true for the thick disc and halo. This similarity is primarily attributed to these components covering similar metallicity regimes, which significantly influences planet occurrence rates in our model. A list of the occurrence rates can be found in Table A.1.
@@
The thick disc and halo bear a distinct planet population compared to the more metal-rich regions. The lower metallicity of the stars favour low-mass planet formation over more massive ones. The occurrence rate of Earth-like planets is a 1.5 times larger than for the bulge and thin disc component, at 5.5 planets per star, with a comparable drop in the occurrence rate of super-Earths to 2.5 planets per star. More striking is the difference in occurrence rates for Neptunian and giant planets. Neptunian planets are five times more common in the thin disc and bulge compared to the thick disc and halo, with frequencies around 0.6. This ratio increases with increasing planet mass, the occurrence rates of giant planets with masses are 20 times higher in the metal-rich components compared to the thick disc and halo. The planets in these regions are also older on average, at 7.7 Gyr for the thick disc and 8.2 Gyr for the halo (see also Table 5).
@@
Taking a higher level view, we analysed the distribution of planet types across the whole galaxy. We show radial profiles in Figure 7, while 2D distribution maps are shown in Figure 8. Giant planets have a higher concentration in the centre of the galaxy, correlating with the region’s higher metallicity. Figure 7 shows the normalised planet density as a function of distance from the galactic centre. The spatial density of planets depends on the stellar density and planet occurrence rates. Thus, unsurprisingly, the highest density of planets is found towards the galactic centre due to the large stellar density in this region. Moving away from the galactic centre, the number of giant planets drops faster than for lower mass planets, due to the higher metallicity-dependence of the giant planet occurrence rates. In total, 90% of giant planets and 84% of Earth-like planets can be found within 8 kpc from the galactic centre, roughly the distance of the Sun from the MW centre.
| Bulge | Thin disc | Thick disc | Halo | |
| Earth-Like | 8.0 | 6.1 | 7.8 | 8.3 |
| Super-Earth | 7.5 | 5.7 | 7.5 | 8.0 |
| Neptunian | 6.8 | 5.1 | 7.2 | 7.3 |
| Sub-giant | 6.9 | 5.1 | 6.9 | 6.6 |
| Giant | 6.5 | 4.8 | 6.8 | 6.1 |
@@
Further, we find that the total number of planets is dependent on the assumed number of initial planet embryos. While the dependence is relatively weak for massive planets – echoing the conclusions of Emsenhuber et al. (2021a) – a more pronounced dependence is observed for lower mass planets. We calculated the total number of planets within kpc of the HESTIA galaxy for = 10, 20, 50, and 100 and tabulate them in Table 6. Depending on the assumed number of embryos, this yields between and planets around Sun-like stars in the HESTIA analogue. It is probable that the number of embryos is not universal for protoplanetary systems and depends, amongst other things on the metallicity (see Section 5.3.2). The = 10 and = 100 cases can however be treated as a lower and upper estimate for the total number of planets.
| 10 | 20 | 50 | 100 | |
| Earth | 3.05 | 5.18 | 9.36 | 11.00 |
| Super-Earth | 3.73 | 5.14 | 9.09 | 10.97 |
| Neptunian | 1.25 | 1.37 | 1.31 | 1.42 |
| Sub-giant | 0.35 | 0.38 | 0.29 | 0.32 |
| Giant | 0.82 | 0.82 | 0.71 | 0.74 |
| Total | 9.21 | 12.89 | 20.77 | 24.45 |
@@ 4.2 Low-metallicity environments
To quantify this effect, we considered an alternative metallicity model, where we assigned zero planets to stars below the metallicity threshold [Fe/H] = -0.6. We compared the occurrence rates to the previous prescriptions. Figure 9 shows the ratio between the occurrence rate for the fiducial and this truncated model for the different planet types and galaxy components. The variations are minimal for the bulge and thin disc, since most stars in these regions are above the metallicity threshold. For the low-metallicity environments, the occurrence rates of low-mass planets in the halo is up to 50% lower when assuming this more restrictive metallicity assumption.
Interestingly, the occurrence rate for giant planets does not significantly differ between the two models, not even in the low-metallicity galaxy components such as the halo, even though the occurrence rates of these planets have the highest metallicity dependence. This stems from the fact that, within our model, giant planets already show near-zero occurrence rates at a metallicity of [Fe/H] = -0.6. Consequently, the variation in the metallicity models does not significantly change their estimated frequency. In contrast, the occurrence rates for Earth-like planets show severe reduction. Given that the occurrence of lower mass planets is only weakly dependent on metallicity, and considering that approximately half of the star particles in the halo have metallicities below our cutoff, the occurrence rate for Earth-like planets is effectively halved in this scenario.
It is important to note that the actual MW halo is more metal-poor than that of the HESTIA analogue. The effect demonstrated here is therefore likely to be even more pronounced in reality.
@@ 4.3 Planets around intermediate- and low-mass dwarfs
For lower mass stars, the correlation between planet occurrence rates with metallicity becomes weaker, although the average metallicity to form a given planet type is larger (Burn et al., 2021). Consequently, the planet populations in the different regions of the galaxy are more homogeneous compared to those around Sun-like stars. The occurrence rates are shown in Figure 10 and Table A.1.
Around intermediate-mass dwarfs (), giant planets are extremely rare in the thick disc and halo with occurrence rates of 0.001 – 0.002. These values increase by a factor of 10 for the metal-rich regions. The occurrence rate of Earth-like planets for these stars is rather insensitive to the metallicity and and stays between four and five planets per star. Super-Earths on the other hand are more metallicity-dependent with occurrence rates around one planet per star for the metal poor region, three times lower than for the metal rich regions.
@@
The weaker metallicity dependence also affects the trends in the occurrence rates with stellar mass (see Table A.1). For massive planets (Neptunians, sub-giants, giants), the occurrence rates consistently increase with stellar mass in all components. In contrast, the occurrence rate of Earth-like planets decreases with stellar mass in the metal-rich regions (bulge, thin disc) but increases in the metal-poor regions (thick disc, halo). For Super-Earths, the occurrence rates have a minimum at in the metal-poor regions.
@@ 5.1 Planet populations in the galactic context
Our results show that galactic chemical evolution and variations in stellar populations profoundly affect planet demographics throughout the galaxy. For Sun-like stars, we observe that metal-rich regions, specifically the bulge and thin disc of the MW, harbour the largest number of planets with masses , in both absolute numbers and relative frequencies. In contrast, the metal-poor regions, namely, the thick disc and halo, show a significant suppression in giant planet formation, which is attributed to the lack of solid material building blocks, which are crucial for forming the massive planet cores before the dispersal of the protoplanetary gas discs. Specifically, the occurrence rate of giant planets around Sun-like stars is an order of magnitude higher in the metal-rich regions. Earth-like planets and Super-Earths are more insensitive to the metallicity variations, but still affected. Earth-like planets around Sun-like stars are actually most frequent in the low metallicity environments of the thick disc and halo, being around twice as common as in the thin disc. Super-Earths show the opposite trend and are about half as frequent in the low metallicity regions.
For lower mass stars, with and , we observe that planet populations are less dependent on metallicity, leading to a more uniform distribution across the galaxy. The lower amounts of solid material available in the protoplanetary discs around these stars generally results in host stars having a higher average metallicity compared to the Sun-like case for the same planet type. Thus, giant and sub-giant planets remain significantly rarer in low-metallicity regions. No giant planets with masses greater are found around any stars of masses , regardless of the galactic environment.
Occasionally, the concept of a galactic habitable zone is discussed in the literature (Gonzalez et al., 2001; Lineweaver, 2001; Prantzos, 2008; Gowanlock et al., 2011), suggesting that life-supporting conditions are more likely in certain galactic regions, including some of our previous work involving hydrodynamical galaxy simulations (Forgan et al., 2017). The key idea is that the galactic centre, with its extreme environmental conditions, is less conducive to life due to higher radiation levels and more frequent stellar encounters. Conversely, the outer regions are thought to be too metal-depleted for habitable planet formation. While we have not explored the radiation-related arguments in the inner region, we find no evidence of an outer limit. Low-mass, rocky planets remain abundant throughout the galaxy far into the stellar halo. Moreover, some of these models suggest that giant planets might be detrimental to habitability or the formation of Earth-like planets (Lineweaver, 2001; Forgan et al., 2017). Under this assumption, stars in the far-out metal-poor regions could be considered favourable, since giant planet formation is suppressed in these regions.
@@ 5.2 Empirical occurrence rates
The robustness of our results applied to the real MW depends on the accuracy of the NGPPS occurrence rates compared to true occurrence rates. Matuszewski et al. (2023) conducted a planet yield analysis for the PLATO mission (Rauer et al., 2014), and compared the occurrence rate estimates around FGK stars from various sources. This comparison includes the NGPPS run, and two empirical models based on Kepler data - one by Hsu et al. (2019) and the other by Kunimoto & Matthews (2020). Unlike the mass-based categorization in NGPPS, these models focus on planet radius, aligning with the observational constraints of transit missions.
@@ 5.3.1 Disc lifetimes
The lifetime of protoplanetary discs, primarily governed by external photoevaporation, plays a crucial role in planet formation. We adopt the range of photoevaporation rates from the Emsenhuber et al. (2021b) NGPPS sample, which are chosen to reproduce typical disc lifetimes in the solar neighbourhood. Efficient massive planet formation () in the Bern model require disc lifetimes around 3 – 6 Myr, indicating that early disc dispersal is not a dominant mechanism with respect to halting giant formation (Schlecker et al., 2021). This is reflected in the weak correlation coefficient between photoevaporation rate and planet type (Figure 2).
However, it is important to note that most studies of disc lifetimes, including those by Emsenhuber et al. (2021a) and Ansdell et al. (2017), have focussed on regions with relatively low stellar densities. In contrast, the galactic centre presents a drastically different environment, where higher stellar densities lead to more intense photoevaporation and, consequently, shorter disc lifetimes (Winter & Haworth, 2022). The Central Molecular Zone (CMZ) in particular, exhibits gas surface density of (Henshaw et al., 2016), exceeding that of the solar neighbourhood by a factor of 100 (Kruijssen, 2012), which is connected to larger and denser stellar clusters (Farias et al., 2023). This heightened density can significantly amplify photoevaporation, and thus reduce disc lifetime. Winter et al. (2020) constructed an analytical model for the disc lifetime as a function of galactic environment, and find that virtually all discs in the CMZ have lifetimes ¡3 Myr, severely hindering giant planet formation.
While the CMZ is an exceptional region within the MW, with a diameter of approximately 0.6 kpc, it is not sufficiently resolved in our simulations. Nonetheless, it is plausible that the shortened disc lifetimes in high-density regions towards the galactic centre could lower the occurrence rates of giant planets.
@@ 5.3.2 Initial embryo distribution
The formation of planet embryos is not explicitly modeled in the Bern model. Instead, a fixed number of embryos are distributed logarithmically at the beginning of each run. However, an analysis of the connection between initial parameter and resulting planet population for the NGPPS population performed by Schlecker et al. (2021) found the location of a planet embryo is the single strongest predictor for the final planetary type and this distribution might vary significantly across different galactic environments.
@@
In terms of the number of planet embryos, they find that the total number of embryos increases with increasing planetesimal disc mass and density slope. For low solid disc masses, specifically between 0.5 and 5 AU, around 20–40 embryos are formed, depending on the assumed density profile. At a higher planetesimal disc mass of , around 60 embryos are formed, irrespective of the assumed profile. Translated into the context of our planet population model, which assumes universal distribution of gas disc masses the galaxy, these findings imply that stars in metal-rich regions (boasting higher solid disc masses) would, on average, form a greater number of planet embryos. Since the occurrence rates of low-mass planets in the model increases with the initial number of embryos (Table 6), this would indicate that Earth-like planets and Super-Earths are more common in metal-rich regions. The frequencies of giant planets on the other hand are insensitive to and are not strongly affected.
5.4 Relation to previous works
@@
Bashi & Zucker (2022) analysed a sample of 506 planet candidates orbiting around 369 FGK host stars, categorising them into thin- and thick-disc components. They found that the occurrence rate of close-in super-Earths in the thin disc is approximately 1.3-3.6 larger than in the thick disc, albeit with low number statistics as only 8 of their planets are associated with thick disc stars. This is consistent with our findings, where the lower average [Fe/H] in the thick disc leads to an increased occurrence rate of Earth-like planets, with a corresponding drop in the frequency of super-Earths by a factor of 1.5 . Bashi & Zucker (2022), however, attribute this variation primarily to the dynamical history of the stars, including potential disruptive encounters, and gravitational perturbations among planets in typically older systems, rather than differences in metallicity.